The Siler Construction Company is about to bid on a new industrial construction project. To formulate their bid, the company needs to estimate the time required for the project. Based on past experience, management expects that the project will require at least 24 months, and could take as long as 48 months if there are complications. The most likely scenario is that the project will require 30 months.

a) Assume that the actual time for the project can be approximated using a triangular probability distribution. What is the probability that the project will take less than 30 months? If required, round your answer to two decimal places.

b) What is the probability that the project will take between 28 and 32 months? If required, round your answer to four decimal places.
Answer #1

for traingular distribution:

parameter:

a= 24
b= 48
m = 30

a)

P(X<30 ) =(x-a)2/((b-a)*(m-a)) =(30-24)^2/((48-24)*(30-24))= 0.2500

b)

P(28<X<32)=P(X<32)-P(X<28) =0.4074-0.1111 =0.2963